OpenStudy (anonymous):

someone teach me algebra in 2 seconds

OpenStudy (anonymous):

its math :)

OpenStudy (anonymous):

I wish that was possible

OpenStudy (anonymous):

no can do.

OpenStudy (anonymous):

\[m=\frac{y_2-y_1}{x_2-x_1}\] \[ax^2+bx+c=0\iff x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

quadratic formula right? @satellite73

OpenStudy (anonymous):

2 seconds passed :/ we failed :/ lol :)

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

 quadratic formula slope

OpenStudy (anonymous):

Lol @someone

OpenStudy (anonymous):

LolQ what a lovely face!

OpenStudy (anonymous):

@help123please. lol that is not be.. the person who you tagged... hahaha

OpenStudy (anonymous):

*not me

OpenStudy (anonymous):

who has taken the algebra end of course

OpenStudy (anonymous):

for what?

OpenStudy (anonymous):

I know 'tis not you.

OpenStudy (anonymous):

I just didn't feel like typing 'someone' twice. XD

OpenStudy (anonymous):

laziness:P

OpenStudy (anonymous):

lol its @some_someone hahah xD

OpenStudy (anonymous):

Indeed,chooch.

OpenStudy (anonymous):

I wish we could drink in information.

OpenStudy (anonymous):

lol yeah

OpenStudy (anonymous):

Blend the information needed for each class into a milkshake kind of thing and drink it.

OpenStudy (anonymous):

what

OpenStudy (anonymous):

yep :P

OpenStudy (anonymous):

K,I shall keep calm and call you Batman,Batman.

OpenStudy (anonymous):

haha xP *putting on the suit* c; *getting the BATMOBILE ready* cx lol on my way -----> *flying away* hahaha

OpenStudy (anonymous):

XDDDDD.

OpenStudy (anonymous):